# EW D-optimal Designs for Experiments with Mixed Factors

### v1

@@ 5 Conclusion
Proof of Theorem S3.2.
According to Theorem 2.1, a necessary and sufficient condition for a design $\boldsymbol{\xi}$ to be EW D-optimal is $\max_{\mathbf{x}\in\mathcal{X}}d(\mathbf{x},\boldsymbol{\xi})\leq p$, which is equivalent to $\max_{\mathbf{x}\in\mathcal{X}}{\rm tr}({\mathbf{F}}(\boldsymbol{\xi})^{-1}{\mathbf{F}}_{\mathbf{x}})\leq p$. Then the conclusions are obtained due to (S.8), (S.9), and ${\mathbf{F}}(\boldsymbol{\xi})=\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}$ ∎
S5 Model Selection and Design Comparison for Paper Feeder Experiment
@@
S5.1 Model selection for paper feeder experiment
For the paper feeder experiment, the summarized response variables include the number of misfeeds $y_{i1}$ , the number of precise feeding of the paper $y_{i2}$ , and the number of multifeeds $y_{i3}$ at the experimental setting ${\mathbf{x}}_{i}$ It is important to notice that these two types of failures, misfeed or multifeed, cannot occur simultaneously. Therefore, a multinomial logistic model is more appropriate than two separate generalized linear models for analyzing this experiment. For those three-level factors, namely $x_{2}$, $x_{3}$, $x_{5}$, $x_{6}$, $x_{7}$, and $x_{8}$ which are all quantitative factors, we follow Wu and Hamada (2009) and Joseph and Wu (2004) to convert the two degrees of freedom for each factor into linear and quadratic components with contrasts $(-1,0,1)$ and $(1,-2,1)$, respectively. Following Joseph and Wu (2004), we code each of the two-level factors $x_{1}$ and $x_{4}$ , which are qualitative factors, as $(-1,1)$. As for the continuous factor stack force $M$, we transform it to $\log(M+1)$, which is slightly different from $\log M$ in Joseph and Wu (2004), to deal with a few cases with $M=0$.
@@
$\log\frac{\pi_{i1}}{\pi_{i2}+\pi_{i3}}$ $=$ $\beta_{11}+\beta_{12}\log(M_{i}+1)+\beta_{13}x_{i1}+\beta_{14}x_{i2l}+\beta_{15}x_{i2q}+\beta_{16}x_{i3l}+\beta_{17}x_{i3q}$
$+$ $\beta_{18}x_{i4}+\beta_{19}x_{i5l}+\beta_{110}x_{i5q}+\beta_{111}x_{i6l}+\beta_{112}x_{i6q}+\beta_{113}x_{i7l}+\beta_{114}x_{i7q}$
$+$ $\beta_{115}x_{i8l}+\beta_{116}x_{i8q}\ ,$
$\log\frac{\pi_{i1}+\pi_{i2}}{\pi_{i3}}$ $=$ $\beta_{21}+\beta_{22}\log(M_{i}+1)+\beta_{23}x_{i1}+\beta_{24}x_{i2l}+\beta_{25}x_{i2q}+\beta_{26}x_{i3l}+\beta_{27}x_{i3q}$
$+$ $\beta_{28}x_{i4}+\beta_{29}x_{i5l}+\beta_{210}x_{i5q}+\beta_{211}x_{i6l}+\beta_{212}x_{i6q}+\beta_{213}x_{i7l}+\beta_{214}x_{i7q}$
$+$ $\beta_{215}x_{i8l}+\beta_{216}x_{i8q}\ ,$
@@
S5.2 Locally D-optimal designs and rounding algorithm for paper feeder experiment
@@
$\boldsymbol{\hat{\boldsymbol{\theta}}}=$ $(\hat{\beta}_{11},\hat{\beta}_{12},\hat{\beta}_{13},\hat{\beta}_{14},\hat{\beta}_{15},\hat{\beta}_{16},\hat{\beta}_{17},\hat{\beta}_{18},\hat{\beta}_{19},\hat{\beta}_{110},\hat{\beta}_{111},\hat{\beta}_{112},\hat{\beta}_{113},\hat{\beta}_{114},\hat{\beta}_{115},\hat{\beta}_{116},$
$\hat{\beta}_{21},\hat{\beta}_{22},\hat{\beta}_{23},\hat{\beta}_{24},\hat{\beta}_{25},\hat{\beta}_{26},\hat{\beta}_{27},\hat{\beta}_{28},\hat{\beta}_{29},\hat{\beta}_{210},\hat{\beta}_{211},\hat{\beta}_{212},\hat{\beta}_{213},\hat{\beta}_{214},\hat{\beta}_{215},\hat{\beta}_{216})^{T}$
$=$ (7.995,3.268,1.275,1.531,0.044,0.156,0.141,0.534,0.261,0.418(7.995,-3.268,-1.275,1.531,0.044,-0.156,-0.141,0.534,-0.261,0.418,( 7.995 , - 3.268 , - 1.275 , 1.531 , 0.044 , - 0.156 , - 0.141 , 0.534 , - 0.261 , 0.418 ,
$-1.749,-0.084,-0.207,0.759,0.782,0.356,10.928,-2.461,-0.409,0.711,$
$0.080,-0.144,-0.120,0.196,0.019,-0.023,-0.931,-0.012,-0.133,0.153,$
$0.128,0.125)^{T}\ .$
@@
The ForLion exact designs (with grid-0.1, grid-0.5, grid-1, and grid-2.5) in Table S.2 also show the convenience by adopting our rounding algorithm (Algorithm 2). Once an optimal approximate design is obtained, the exact design with a user-specified grid level can be obtained instantly with high relative efficiency. To show how we choose the merging threshold $\delta_{r}=3.2$ for this experiment, in Figure S.1 we show how the relative efficiency (left panel) and the number of design points (right panel) change along with the relative distance (i.e., the merging threshold $\delta_{r}$), where the grid levels 0.1, 0.5, 1 and 2.5, and the relative distances between 2 and 6 with increment of 0.1 are used for illustration. Notably, the next change on the number of design points after $\delta_{r}=3.2$ will not occur until $\delta_{r}=31.29$, which is too large. As a result, we recommend the merging threshold $\delta_{r}=3.2$ for this experiment. The choice of grid level typically depends control accuracy of the factors and the options that the experimenter may have. As illustrated in Table S.2, the experimenter may choose the smallest possible grid level if applicable.
@@
S5.3 More graphs and tables for EW D-optimal designs for paper feeder experiment
@@ 18 Original 10 5 10 10 5 5 10 5 10 5 10 10 5
Table S.4: Designs constructed on $[0,160]$ of stack force for paper feeder experiment
@@
S6 Robustness of Sample-based EW Designs
@@
For illustration purpose, we simulate five random sets of parameter vectors from the prior distributions listed in Table S1 of Huang et al. (2024), denoted by $\boldsymbol{\Theta}^{(k)}=\{\boldsymbol{\theta}_{1}^{(k)},\ldots,$ $\boldsymbol{\theta}_{B}^{(k)}\}$, with $B=1000$, and $k\in\{1,2,3,4,5\}$. For each $k=1,\ldots,5$, we obtain the corresponding sample-based EW D-optimal design $\boldsymbol{\xi}^{(k)}$ for the minimizing surface defects experiment, which maximizes $f_{\rm SEW}(\boldsymbol{\xi})=|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta}^{(k)})\}|=|B^{-1}\sum_{j=1}^{B}{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta}_{j}^{(k)})|$. The numbers of design points are 15, 13, 21, 19, and 17, respectively.
For $l,k\in\{1,2,3,4,5\}$, we calculate the relative efficiencies of $\boldsymbol{\xi}^{(l)}$ with respect to $\boldsymbol{\xi}^{(k)}$ under $\boldsymbol{\Theta}^{(k)}$, that is,
$\left(\frac{|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}^{(l)},\boldsymbol{\Theta}^{(k)})\}|}{|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}^{(k)},\boldsymbol{\Theta}^{(k)})\}|}\right)^{1/p}$
@@
S7.1 Electrostatic discharge experiment
@@
Table S.5: Factors and parameters for electrostatic discharge (ESD) experiment
@@
Table S.6: EW D-optimal design (left) and a corresponding exact design (right) for electrostatic discharge (ESD) experiment
@@
To make comparison, we randomly generate $N=10,000$ parameter vectors $\{\boldsymbol{\theta}_{1},\ldots,$ $\boldsymbol{\theta}_{N}\}$ from the prior distribution listed in Table S.5. For each $j\in\{1,\ldots,N\}$ and each $\boldsymbol{\xi}$ under comparison, we compute ($|{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta}_{j})|^{1/p}=|{\mathbf{X}}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\theta}_{j}){\mathbf{X}}_{\boldsymbol{\xi}}|^{1/p}$ with $p=7$ in this case. In Figure S.3, we present the resulting values as frequency polygons for our EW ForLion design, the locally D-optimal design (“ForLion”) obtained by Huang et al. (2024), and the $d$-QPSO design (“PSO”) obtained by Lukemire et al. (2019). Compared with ForLion design or PSO design, the frequency polygon for EW ForLion design shows a slightly higher density with low objective functions values, but also a noticeably increased frequency in the higher-value range. Actually, in Table S.7, it clearly shows that the overall mean and median objective function values for EW ForLion designs (approximate and exact designs listed in Table S.6) are higher than ForLion and PSO designs.
@@
Table S.7: Mean and median objective function values based on 10,000 simulated parameter vectors for electrostatic discharge (ESD) experiment
@@
S7.2 A three-continuous-factor example under a GLM
This example refers to Example 4.7 of Stufken and Yang (2012) on a logistic regression model with three continuous factors, represented as ${\rm logit}(\mu_{i})=\beta_{0}+\beta_{1}x_{i1}+\beta_{2}x_{i2}+\beta_{3}x_{i3}$ , where the factors were originally defined as $x_{i1}\in[-2,2]$, $x_{i2}\in[-1,1]$ and $x_{i3}\in(-\infty,\infty)$.

### v2

@@ 5 Conclusion
Proof of Theorem 6.
According to Theorem 2, a necessary and sufficient condition for a design $\boldsymbol{\xi}$ to be EW D-optimal is $\max_{\mathbf{x}\in\mathcal{X}}d(\mathbf{x},\boldsymbol{\xi})\leq p$, which is equivalent to $\max_{\mathbf{x}\in\mathcal{X}}{\rm tr}({\mathbf{F}}(\boldsymbol{\xi})^{-1}{\mathbf{F}}_{\mathbf{x}})\leq p$. Then the conclusions are obtained due to ${\mathbf{F}}(\boldsymbol{\xi})=\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}$ ∎
S4 Model Selection and Design Comparison for Paper Feeder Experiment
@@
S4.1 Model selection for paper feeder experiment
For the paper feeder experiment, the summarized response variables include the number of misfeeds $y_{i1}$ , the number of precise feeding of the paper $y_{i2}$ , and the number of multifeeds $y_{i3}$ at the experimental setting ${\mathbf{x}}_{i}$ It is important to notice that these two types of failures, misfeed or multifeed, cannot occur simultaneously. Therefore, a multinomial logistic model is more appropriate than two separate generalized linear models for analyzing this experiment. For those three-level factors, namely $x_{2}$, $x_{3}$, $x_{5}$, $x_{6}$, $x_{7}$, and $x_{8}$, which are all quantitative factors, we follow Wu and Hamada (2009) and Joseph and Wu (2004) to convert the two degrees of freedom for each factor into linear and quadratic components with contrasts $(-1,0,1)$ and $(1,-2,1)$, respectively. Following Joseph and Wu (2004), we code each of the two-level factors $x_{1}$ and $x_{4}$ , which are qualitative factors, as $(-1,1)$. As for the continuous factor stack force $M$, we transform it to $\log(M+1)$, which is slightly different from $\log M$ in Joseph and Wu (2004), to deal with a few cases with $M=0$.
@@
$\log\frac{\pi_{i1}}{\pi_{i2}+\pi_{i3}}$ $=$ $\beta_{11}+\beta_{12}\log(M_{i}+1)+\beta_{13}x_{i1}+\beta_{14}x_{i2l}+\beta_{15}x_{i2q}$
$+$ $\beta_{16}x_{i3l}+\beta_{17}x_{i3q}+\beta_{18}x_{i4}+\beta_{19}x_{i5l}+\beta_{110}x_{i5q}+\beta_{111}x_{i6l}$
$+$ $\beta_{112}x_{i6q}+\beta_{113}x_{i7l}+\beta_{114}x_{i7q}+\beta_{115}x_{i8l}+\beta_{116}x_{i8q}\ ,$
$\log\frac{\pi_{i1}+\pi_{i2}}{\pi_{i3}}$ $=$ $\beta_{21}+\beta_{22}\log(M_{i}+1)+\beta_{23}x_{i1}+\beta_{24}x_{i2l}+\beta_{25}x_{i2q}$
$+$ $\beta_{26}x_{i3l}+\beta_{27}x_{i3q}+\beta_{28}x_{i4}+\beta_{29}x_{i5l}+\beta_{210}x_{i5q}+\beta_{211}x_{i6l}$
$+$ $\beta_{212}x_{i6q}+\beta_{213}x_{i7l}+\beta_{214}x_{i7q}+\beta_{215}x_{i8l}+\beta_{216}x_{i8q}\ ,$
@@
S4.2 Locally D-optimal designs for paper feeder experiment
@@
$\hat{\boldsymbol{\theta}}$ $=$ $(\hat{\beta}_{11},\hat{\beta}_{12},\hat{\beta}_{13},\hat{\beta}_{14},\hat{\beta}_{15},\hat{\beta}_{16},\hat{\beta}_{17},\hat{\beta}_{18},$
$\hat{\beta}_{19},\hat{\beta}_{110},\hat{\beta}_{111},\hat{\beta}_{112},\hat{\beta}_{113},\hat{\beta}_{114},\hat{\beta}_{115},\hat{\beta}_{116},$
$\hat{\beta}_{21},\hat{\beta}_{22},\hat{\beta}_{23},\hat{\beta}_{24},\hat{\beta}_{25},\hat{\beta}_{26},\hat{\beta}_{27},\hat{\beta}_{28},$
$\hat{\beta}_{29},\hat{\beta}_{210},\hat{\beta}_{211},\hat{\beta}_{212},\hat{\beta}_{213},\hat{\beta}_{214},\hat{\beta}_{215},\hat{\beta}_{216})^{T}$
$=$ (7.995,3.268,1.275,1.531,0.044,0.156,0.141,0.534(7.995,-3.268,-1.275,1.531,0.044,-0.156,-0.141,0.534,( 7.995 , - 3.268 , - 1.275 , 1.531 , 0.044 , - 0.156 , - 0.141 , 0.534 ,
$-0.261,0.418,-1.749,-0.084,-0.207,0.759,0.782,0.356,$
$10.928,-2.461,-0.409,0.711,0.080,-0.144,-0.120,0.196,$
$0.019,-0.023,-0.931,-0.012,-0.133,0.153,0.128,0.125)^{T}\ .$
@@
The ForLion exact designs (with grid-0.1, grid-0.5, grid-1, and grid-2.5) in Table S.2 also show the convenience by adopting our rounding algorithm (Algorithm 2). Once an optimal approximate design is obtained, the exact design with a user-specified grid level can be obtained instantly with high relative efficiency. To show how we choose the merging threshold $\delta_{r}=3.2$ for this experiment, in Figure S.1 we show how the relative efficiency (left panel) and the number of design points (right panel) change along with the relative distance (i.e., the merging threshold $\delta_{r}$), where the grid levels 0.1, 0.5, 1 and 2.5, and the relative distances between 2 and 6 with increment of 0.1 are used for illustration. Notably, the next change on the number of design points after $\delta_{r}=3.2$ will not occur until $\delta_{r}=31.29$, which is too large. As a result, we recommend the merging threshold $\delta_{r}=3.2$ for this experiment. The choice of grid level typically depends control accuracy of the factors and the options that the experimenter may have. As illustrated in Table S.2, the experimenter may choose the smallest possible grid level when applicable.
@@
S4.3 More graphs and tables for paper feeder experiment
@@ 18 Original 10 5 10 10 5 5 10 5 10 5 10 10 5
Table S.4: Designs on $[0,160]$ of stack force for paper feeder experiment
@@
S5 Robustness of Sample-based EW Designs
@@
For illustration purpose, we simulate five random sets of parameter vectors from the prior distributions listed in Table S1 of Huang et al. (2024), denoted by $\boldsymbol{\Theta}^{(r)}=\{\boldsymbol{\theta}_{1}^{(r)},\ldots,$ $\boldsymbol{\theta}_{B}^{(r)}\}$, with $B=1000$, and $r\in\{1,2,3,4,5\}$. For each $r=1,\ldots,5$, we obtain the corresponding sample-based EW D-optimal design $\boldsymbol{\xi}^{(r)}$ for the minimizing surface defects experiment, which maximizes $f_{\rm SEW}(\boldsymbol{\xi})=|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta}^{(r)})\}|=|B^{-1}\sum_{j=1}^{B}{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta}_{j}^{(r)})|$. The numbers of design points are 15, 13, 21, 19, and 17, respectively.
For $l,r\in\{1,2,3,4,5\}$, we calculate the relative efficiencies of $\boldsymbol{\xi}^{(l)}$ with respect to $\boldsymbol{\xi}^{(r)}$ under $\boldsymbol{\Theta}^{(r)}$, that is,
$\left(\frac{|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}^{(l)},\boldsymbol{\Theta}^{(r)})\}|}{|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}^{(r)},\boldsymbol{\Theta}^{(r)})\}|}\right)^{1/p}$
@@
S6.1 Electrostatic discharge experiment
@@
Table S.5: Factors and parameters for electrostatic discharge experiment
@@
Table S.6: EW D-optimal approximate design (left) and exact design (right, $n=100$) for electrostatic discharge experiment
@@
To make comparison, we randomly generate $N=10,000$ parameter vectors $\{\boldsymbol{\theta}_{1},\ldots,$ $\boldsymbol{\theta}_{N}\}$ from the prior distribution listed in Table S.5. For each $j\in\{1,\ldots,N\}$ and each $\boldsymbol{\xi}$ under comparison, we compute ($|{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta}_{j})|^{1/p}=|{\mathbf{X}}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\theta}_{j}){\mathbf{X}}_{\boldsymbol{\xi}}|^{1/p}$ with $p=7$ in this case. In Figure S.3, we present the resulting values as frequency polygons for our EW ForLion design, the locally D-optimal design (“ForLion”) obtained by Huang et al. (2024), and the $d$-QPSO design (“PSO”) obtained by Lukemire et al. (2019). Compared with the ForLion or PSO design, the frequency polygon for EW ForLion design shows a slightly higher density with low objective functions values, but also a noticeably increased frequency in the higher-value range. Actually, in Table S.7, it clearly shows that the overall mean and median objective function values for EW ForLion designs (approximate and exact designs listed in Table S.6) are higher than ForLion and PSO designs.
@@
Table S.7: Mean and median objective function values based on 10,000 simulated parameter vectors for electrostatic discharge experiment
@@
S6.2 A three-continuous-factor example under a GLM
This example refers to Example 4.7 of Stufken and Yang (2012) on a logistic regression model with three continuous factors, represented as ${\rm logit}(\mu_{i})=\beta_{0}+\beta_{1}x_{i1}+\beta_{2}x_{i2}+\beta_{3}x_{i3}$ , where the factors were originally defined as $x_{i1}\in[-2,2]$, $x_{i2}\in[-1,1]$, and $x_{i3}\in(-\infty,\infty)$.